Give exact and approximate solutions to three decimal places
y^-20y+100=4

Answer by algebrahouse.com(1659)

y² - 20y + 100 = 4
y² - 20y + 96 = 0 {subtracted 4 from both sides}
(y - 12)(y - 8) = 0 {factored into two binomials}
y - 12 = 0 or y - 8 = 0 {set each factor equal to 0}
y = 12 or y = 8 {solved each equation for x}

For more help from me, visit: www.algebrahouse.com